The perimeter fence of a rectangular community park is 310cm. Its length is 10m less than twice its width. Find the dimensions of the park
Answer by macston(5194)

NOTE: I changed the perimeter from 310 cm to 310 m.
W=width; L=length=2W-10m; P=perimeter=2(L+W)=310m
P=2(L+W) Divide each side by 2.
310m/2=L+W Substitute for L.
155m=2W-10m+W Add 10m to each side.
165m=3W Divide each side by 3.
55m=W ANSWER 1: The width is 55 meters.
L=2W-10m=2(55m)-10m=100m ANSWER 2: The length is 100 m.
CHECK:
P=2(L+W)
310 m=2(100m+55M)
310m=2(155m)
310m=310m